OpenStudy (anonymous):

how to solve for p[oH] p[H] on a calculator (i understand how it works but the numbers do not come out correctly)

OpenStudy (mos1635):

it depence on the calc... try: log value =

OpenStudy (anonymous):

hmm i just have to plug in - then log it prints out -log (

OpenStudy (anonymous):

its a TI 30x :/

OpenStudy (mos1635):

lets try something.. type: log 100 =

OpenStudy (mos1635):

what is it say?

OpenStudy (anonymous):

2

OpenStudy (mos1635):

log 0.001 = ?

OpenStudy (anonymous):

-3

OpenStudy (mos1635):

ok

OpenStudy (mos1635):

- log 0.001 = ?

OpenStudy (anonymous):

3

OpenStudy (mos1635):

working good so: - log (value of [H] ) = .....

OpenStudy (anonymous):

thats the part i do not understand OH = 3.0 * 10 ^-9 example H = 3.3 * 10^-6 how did the book get OH ?

OpenStudy (anonymous):

vice versa also :/

OpenStudy (mos1635):

pOH + pH = pKw

OpenStudy (mos1635):

[H+] *[OH-] = Kw

OpenStudy (mos1635):

Kw=?

OpenStudy (mos1635):

lets take it from the top... what you know what to find?

OpenStudy (mos1635):

that is the yheory mumbers...???

OpenStudy (anonymous):

i only got 1.0 * 10^14 = 3.3 * 10^-6 [OH]

OpenStudy (anonymous):

but how do i punch it in on a calc?

OpenStudy (mos1635):

1.0*10^-14 seems like Kw so Kw = [ H+] [ OH-] => [ H+]= Kw / [ OH-] => [ H+]= 1.0 * 10^-14 / 3.3 * 10^-6 => [ H+]= 0.3 * 10^-8 => [ H+]= 3 * 10^-9 so pH=-log[ H+] => pH=-log 3 * 10^-9 pH=8.5

OpenStudy (mos1635):

punch: 1 exp - 14 / 3.3 exp - 6 = log = +-

OpenStudy (mos1635):

is it working???